OpenStudy (anonymous):

The rate of change of N with respect to s is inversely proportional to s. Write and solve the differential equation.

OpenStudy (anonymous):

dN/ds = k/s, where k is a constant.

OpenStudy (anonymous):

How do I find the general solution if there is no N in the equation?

OpenStudy (anonymous):

familiar with separable DE?

OpenStudy (anonymous):

maybe? I will probably remember after you refresh my memory? :)))?

OpenStudy (anonymous):

dN/ds = k/s dN = k (1/s) ds integrate both sides N = k ln|s| + C

OpenStudy (anonymous):

C is also an arbitrary constant

OpenStudy (anonymous):

Oh I had that! I worked that out! I guess I just didn't know what I was doing.